Project

General

Profile

Actions

2016-09-16 physics

Demo paper

  • We have a final round of comments, aim is to finish by end of the week; redistribute to commenters only (it has had two goes around the collaboration)

Emittance paper

  • Victoria has added diffuser cuts; they look quite successful
  • Need to calculate systematic error from "TOF bias" and "Hall probes" and then decide if they are acceptable
    • Scott Wilbur and Victoria/Joe Langlands are chasing respectively
  • Need to chase out MC issue
    • Dipoles (Paolo/Rogers) or target simulation (Ao)?

Scattering paper

  • Ryan is working through systematic errors
    • Looking at the "1/p" relationship and systematic due to momentum mis-reconstruction
    • TOF binning
    • Detector Alignment
    • LiH composition

2016/02 Data analysis

  • still stuttering to get an analysis started for 2016/02 (fields on) data/magnet alignment
  • FC/transfer matrix analysis is hung up because we don't have a reconstruction with the detectors alignment in
  • Ao has renewed his vow to look at the "residuals analysis" - I have also tried to pull in JB; let's see if they emerge. I can look once I finish MOM job

2016/03

  • Optics settings delivered
  • Run with magnets w/c 26th September and week following; proposed "operational plan" to Colin, to be discussed; idea is to do optics scans in the day shift with magnet expert in attendance, one scattering measurement in the evening shift with magnet expert on-call, power down magnets overnight. Not yet confirmed.
  • Propose LiH changeover w/c 10th October. Note user run ends 28th October.

Updated by Rogers, Chris over 6 years ago ยท 2 revisions